OpenStudy (anonymous):

The sum of twice a number and 1 is equal to 3 times the positive square root of the number find the number

OpenStudy (anonymous):

let the number be "x" then the given condition can be mathematically expressed as: \[(2x)+1=3\sqrt{x}\\ \text{squaring both sides...}\\ 4x^2+4x+1=9x\\ 4x^2-5x+1=0\\ 4x^2-4x-x+1=0\\ 4x(x-1)-(x-1)=0\\ (4x-1)(x-1)=0\\ 4x-1=0\qquad{\rm OR}\qquad x-1=0\\ x={1\over4}\qquad{\rm OR}\qquad \boxed{\Large x=1} \] we choose the second since "x" is a number and not a fraction

OpenStudy (anonymous):

ok i see thank you so much!!

OpenStudy (anonymous):

yw